# Prospects of constraining on the polarizations of gravitational waves from binary black holes using space- and ground-based detectors

### v1

@@ V.2 Results with ppE parameters
We evaluate the performance of several typical BBHs using different detectors and various parameters. The results for different detectors and their networks are illustrated in Fig. 10.
@@
Regarding space-based detectors, Taiji, which shares a similar configuration with LISA, surpasses LISA across all metrics, demonstrating superior SNR and reduced parameter uncertainty. Moreover, TianQin’s distinct orbital configuration results in inferior performance compared to LISA and Taiji. Additionally, the significant rise in TianQin’s parameter uncertainty at $\iota=\pi/2$, particularly in $\Delta\alpha_{V}$, leads to the upper limit of the box plot for $\Delta\alpha_{V}$ being substantially higher than its median value in the comprehensive results presented in Fig. 10.
@@
When considering the space-based detector network, significant improvements are observed across all metrics compared to individual detectors. Since both LISA and Taiji utilize heliocentric orbits, their network configuration remains stable, maintaining a consistent angle between the detectors. LISA and the three alternative orbital configurations of Taiji exhibit different angles, which introduce variations in the outcomes. A larger angular separation between the detectors correlates with greater coverage of the sky area with high sensitivity, resulting in differences in parameter uncertainty under similar SNR conditions. Overall, the LISA+TJm configuration achieves the most favorable results, with LISA+TJp surpassing LISA+TJc, especially noticeable in parameters like $\Delta\alpha_{V}$ and $\Delta\Omega$. Moreover, it is evident that the LISA+TJ combination outperforms LISA+TQ, and a network of three detectors surpasses two. For more detailed analysis of space detector networks, refer to Ref. [44].
@@
Single ground-based detectors such as Virgo and KAGRA do not perform well in detecting additional polarization parameters and accurately determining the sky position of the source. This limitation arises because a single detector has only one response angle, and a 10-minute GW signal is relatively short compared to the ground-based detector’s observational period (Earth’s rotation period). In the ground-based detector’s coordinate system, a 10-minute SBBH source travels only $2.5^{\circ}$, whereas a 90-day MBHB source moves nearly $90^{\circ}$ in the space-based detector coordinate system. Consequently, the parameter uncertainty range for ground-based detectors is larger. Furthermore, both LIGO and ET have three detectors, but ET detectors are all coplanar, whereas LIGO’s are not. This situation is similar to the superiority of LISA+TJm over LISA+TJc, where the relative positions of LIGO’s three detectors offer greater advantages compared to ET. Therefore, despite the SNR of ET is one order of magnitude higher than LIGO, their $\Delta\Omega$ values are still very similar.
@@
Regarding the ground-based detector network, currently operational LVK cannot fully compensate for the sensitivity differences between second-generation and third-generation detectors. Compared to ET, LVK exhibits a difference of one to two orders of magnitude in the limitations of ppE parameters. Moreover, due to the varied response angles of LVK, $\Delta\Omega$ values are slightly better than those of ET. The LVK+ET combination is the most comprehensive, significantly enhancing network sensitivity. For instance, $\Delta\alpha_{Q}$ and $\Delta\alpha_{D}$ can approach the sensitivity level of space-based detector networks, with $\Delta\alpha_{V}$ potentially surpassing them. Due to the arm length, the $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ of the ground-based detector network is four to six orders of magnitude higher than that of the space-based detector network, making it impossible to distinguish between breathing and longitudinal modes.
V.3 Results with redshift
@@
In our previous calculations and analysis, we assume fixed redshift: $z=0.01$ for SBBH and $z=1$ for MBHB. Different redshifts also lead to variations in the GW signal, potentially impacting the final results. From Eq. (2), the strain is inversely proportional to the luminosity distance, represented as $h\propto 1/D_{L}\sim 1/z$. Roughly speaking, as redshift increases, the strain decreases. Regarding the ppE parameter $\xi$, the partial derivative of the strain with respect to it, $\partial h/\partial\xi\propto 1/z$. Consequently, the FIM is $(\partial h/\partial\xi|\partial h/\partial\xi)\propto 1/z^{2}$, and thus the uncertainty of the ppE parameter scales as $\Delta\xi\propto z$. In other words, the uncertainty of the ppE parameter is directly proportional to the redshift:
$\Delta\xi(z)=\mathcal{K}z+\Delta\xi(0),$ (21)
@@
where $\mathcal{K}$ is a proportionality constant. We calculate the results for SBBH and MBHB at different redshifts, as shown in Table 2.
@@
$10^{5}$5.01 -8.59 -2.50 -0.82
LISA $10^{6}$4.72 -8.84 -2.42 -0.35
$10^{7}$5.45 -10.13 -2.60 1.62
$10^{5}$5.30 -8.90 -2.79 -1.00
Taiji $10^{6}$4.90 -9.06 -2.52 -0.66
$10^{7}$5.74 -10.42 -2.81 1.36
$10^{5}$4.91 -8.53 -2.25 0.92
TianQin $10^{6}$4.43 -8.68 -1.98 2.32
$10^{7}$5.24 -9.95 -1.93 3.79
$3$1.90 -5.71 0.35 7.93
LIGO $20$2.42 -6.83 -0.18 7.41
$100$2.52 -6.54 -0.15 7.38
$3$1.58 -5.42 2.56 10.09
Virgo $20$2.13 -6.55 2.60 9.59
$100$2.19 -6.21 2.88 9.18
$3$1.44 -5.30 2.91 9.59
KAGRA $20$2.11 -6.57 2.82 8.44
$100$1.89 -5.93 3.26 7.80
$3$2.94 -6.62 -0.85 7.08
ET $20$3.37 -7.78 -1.29 6.62
$100$3.71 -7.81 -1.49 6.30
From Eq. (21), it is clear that the uncertainty of the ppE parameters is proportional to the redshift. Additionally, the rate of change in uncertainty with redshift varies for different ppE parameters. Table 2 illustrates that the influence of redshift on $\Delta\alpha_{Q}$ and $\Delta\alpha_{D}$ is negligible, as most $\mathcal{K}$ values are less than $10^{-4}$. That is because these two parameters are directly related to frequency, and the change in GW amplitude due to redshift has a negligible effect on frequency resolution. Moreover, ground-based detectors exhibit relatively larger $\mathcal{K}$ values compared to space-based detectors.
The $\mathcal{K}$ for $\Delta\alpha_{V}$ is significantly larger compared to the first two parameters. Space-based detectors typically have $\mathcal{K}$ values around $10^{-2}$. In contrast, ground-based detectors like Virgo and KAGRA have $\mathcal{K}$ values around $10^{2}$. LIGO and ET perform better, with $\mathcal{K}<1$ in most cases. Those observations are consistent with the analysis in last section, explaining these differences based on detector performance.
Due to the degeneracy in ground-based detectors, $\Delta\alpha_{B}$ ($\Delta\alpha_{L}$) exhibits significantly large $\mathcal{K}$ values, exceeding $10^{6}$. In contrast, space-based detectors, which break this degeneracy, yield noticeably better results. The relationship between the frequency range of MBHBs and the transfer frequency determines the results in Table 2. Smaller mass MBHBs exhibit lower frequencies closer to the transfer frequency, enhancing the resolution of these two modes and leading to smaller $\mathcal{K}$ values. The results are similar for different detectors: Taiji has the smallest transfer frequency, while TianQin has the largest, corresponding to their respective $\mathcal{K}$ values.
@@ V.4 Multiband observation
Space-based detectors can not only observe MBHBs, but also SBBHs. Among tens of thousands of sources, a small number of SBBHs increase in frequency rapidly and merge into the ground-based detector frequency band within a short period, enabling multiband observations by both space- and ground-based detectors [65]. We calculate FIM for two typical SBBHs observed across multiple frequency bands and compared their results with those observed only by ground-based detectors, as shown in Fig. 11.
@@
Multiband observations provide a significant enhancement to ground-based detectors, particularly in resolving the degeneracy between the breathing and longitudinal modes. The addition of space-based detectors effectively breaks the degeneracy, allowing multiband SBBH observations to limit $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ to levels comparable with space-based detector observations of MBHB. Additionally, multiband observations greatly improve the performance of Virgo and KAGRA in $\Delta\alpha_{V}$, reducing them by nearly three orders of magnitude, approaching the levels achieved by LIGO and ET. Furthermore, the $\Delta\alpha_{V}$ level of the ground-based detector network remains higher than that of the space-based detector network, so the improvement from multiband observations is not as significant.
@@
Regarding $\Delta\alpha_{Q}$ and $\Delta\alpha_{D}$, multiband observations have varying improvements for SBBH depending on the mass. The results for SBBH with $M=20\ \mathrm{M_{\odot}}$ do not show significant improvement, while those for $M=100\ \mathrm{M_{\odot}}$ show some improvement. Due to the high sensitivity of ET, multiband observation does not significantly improve ET and LVK+ET, it significantly enhances the performance of only second-generation detectors. Within the space-based detector frequency band, although both types of SBBH have a duration of one year, their frequency variation ranges differ. Specifically, the frequency variation for $M=20\ \mathrm{M_{\odot}}$ SBBHs is 46 mHz, whereas for $M=100\ \mathrm{M_{\odot}}$ it increases to 78 mHz. Greater frequency variations lead to stronger constraints, thereby affecting the observed improvements. Moreover, longer observation times result in larger frequency variations. While four-year observations generally outperform one-year observations, this improvement is not significant. This is because the further away from the merger, the smaller the frequency change over the same time period. For example, in calculations for SBBH with $M=100\ \mathrm{M_{\odot}}$, the frequency change in one year leading up to 0.1 Hz is 78 mHz, but the change over four years is only 87 mHz, indicating a minimal additional change of 9 mHz over three years. Hence, our choice of a one-year observation duration remains reasonable.
@@
The performance of different space-based detectors significantly affects the results of multi band observations, for example, Taiji is superior to LISA. Moreover, the selected cutoff frequency of SBBHs within the space-based detector band is 0.1 Hz, lower than TianQin’s 0.28 Hz transfer frequency. That disparity renders TianQin less effective than LISA and Taiji in observing breathing and longitudinal modes, resulting in a multiband observation improvement for $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ that is two orders of magnitude weaker compared to the other detectors. Conversely, TianQin exhibits significantly enhanced multiband observations for $\Delta\alpha_{Q}$, $\Delta\alpha_{D}$, and $\Delta\alpha_{V}$, because TianQin’s sensitivity within the selected SBBH frequency band surpasses that of the other detectors, leading to superior results. Overall, the addition of space-based detectors compensates for the degeneracy of ground-based detectors in scalar modes and also improves other aspects, enabling the observation results of SBBH to reach the level of MBHB results.
@@ V.5 Multimessenger observation
Apart from multiband observations using space- and ground-based detectors, multimessenger observations, which include the assistance of electromagnetic (EM) observations, can similarly enhance the performance of GW observations. EM observations offer a unique perspective about BBH sources distinct from GW observations. References [66, 67, 68, 69] have demonstrated that mergers of MBHBs can emit EM radiation from accretion disks, while mergers of SBBHs in active galactic nuclei can produce jets, making both scenarios promising targets for EM detectors. Accurate determination of the source’s sky position enables subsequent EM follow-up observations to search for counterparts. The results of EM observations can then serve as valuable priors to reduce uncertainties in GW observations.
@@
The EM effects produced by BBH mergers are expected to be detected by infrared, optical, and X-ray observatories, and the results from these observatories have different impacts on the enhancement of GW. For simplicity, we consider an ideal scenario where EM observations accurately determine specific parameters, ignoring differences between EM detectors. According to Refs. [63, 44], when considering the improvement from EM observations, the corresponding row and column in the FIM are removed to reduce the uncertainties of other parameters in the GW data. That method assesses the performance enhancement of GW observations under ideal multimessenger conditions. We quantify these enhancements when EM observations perfectly determine redshift, sky position, or inclination angle, comparing them against results from GW observations alone. Notably, although this study only focuse on ppE parameters, multimessenger observations also offer varying degrees of enhancement for time, mass, polarization, and other parameters. For details on enhancements in non-ppE parameters, refer to Ref. [44].
@@
For LISA and Taiji, multimessenger observations demonstrate notable enhancements primarily for MBHBs with $M=10^{7}\ \mathrm{M_{\odot}}$. Determining the redshift or inclination angle through multimessenger observations can reduce $\Delta\alpha_{V}$, $\Delta\alpha_{B}$, and $\Delta\alpha_{L}$ by 15% to 24%. Determining the sky position can reduce these parameters by 36% to 49%. Multimessenger observations yield even more significant improvements for TianQin, enhancing detection capabilities across all three typical MBHB masses. When the redshift, inclination angle, or sky position is accurately determined, $\Delta\alpha_{V}$ can be reduced by 40% to 57%. Furthermore, due to arm length, reductions in $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ are more modest, ranging from 9% to 17%.
@@
For ground-based detectors, the impact of multimessenger observations on $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ is considerably smaller compared to multiband observations, as it does not resolve the degeneracy and thus is not considered further. LIGO shows less significant improvement compared to ET, due to its less diverse response angles. Determining the redshift or inclination angle results in a reduction of $\Delta\alpha_{V}$ by 5% to 7%, with negligible changes observed when determining the sky position. In contrast, ET exhibits more substantial enhancements. When any one of the three parameters is determined, $\Delta\alpha_{V}$ can be reduced by 15% to 43%, with larger reductions for higher-mass SBBHs. Moreover, multimessenger observations yield remarkable improvements for Virgo and KAGRA, achieving reductions in $\Delta\alpha_{V}$ by 97% to 99.4%. This improvement is only one order of magnitude lower than that of LIGO and significantly surpasses the outcomes from GW observations alone.
@@
Multimessenger observations, whether for space- or ground-based detectors, do not improve the other two ppE parameters, $\Delta\alpha_{Q}$ and $\Delta\alpha_{D}$. That is because these two ppE parameters are directly tied to GW frequency, where GW observation precision is highest. Theoretically, multimessenger observations could enhance constraints on all ppE parameters, but in practice, some improvements are too marginal. Therefore, we only present results where the parameter uncertainty is reduced by more than 5%, with most unreported results being less than 1%. In summary, multimessenger observations can moderately enhance the performance of GW observations, providing stronger constraints on ppE parameters.

### v2

@@ V.2 Results with ppE parameters
The performance of several typical BBHs is assessed using different detectors and various parameters. The results for different detectors and their networks are depicted in Fig. 10.
@@
Regarding space-based detectors, Taiji, which shares a similar configuration with LISA, surpasses LISA across all aspects, demonstrating superior SNR and reduced parameter uncertainty. TianQin’s unique orbital configuration results in inferior performance compared with LISA and Taiji. Additionally, the significant increase in TianQin’s parameter uncertainty at $\iota=\pi/2$, particularly in $\Delta\alpha_{V}$, causes the upper limit of the box plot for $\Delta\alpha_{V}$ to be substantially higher than its median value as shown in the comprehensive results presented in Fig. 10.
@@
When considering the space-based detector network, significant improvements are observed across all metrics compared with individual detectors. Due to the heliocentric orbits for both LISA and Taiji, their network configuration remains stable, ensuring a consistent angle between the detectors. The three alternative orbital configurations of Taiji, along with LISA, exhibit distinct angles that introduce outcome variations. A larger angular separation between the detectors increases sky area coverage with high sensitivity, resulting in differences in parameter uncertainty under similar SNR conditions. The LISA+TJm configuration demonstrates superior performance overall, with LISA+TJp outperforming LISA+TJc, particularly in parameters such as $\Delta\alpha_{V}$ and $\Delta\Omega$. Furthermore, it is evident that LISA+TJ+TQ surpasses LISA+TJ, highlighting the advantage of a network comprising three detectors over two. For a more comprehensive analysis of space detector networks, please refer to Ref. [47].
@@
The ability of ground-based detectors like Virgo and KAGRA to detect additional polarization and determine the sky position is limited. This limitation stems from the fact that a single detector only has only one response angle, and a 10-minute GW signal is very short compared with the observational period (Earth’s rotation period). In the ground-based detector’s coordinate system, a 10-minute SBBH source travels only $2.5^{\circ}$, whereas a 90-day MBHB source moves nearly $90^{\circ}$ in the space-based detector coordinate system. Consequently, the parameter uncertainty range for ground-based detectors is larger. Furthermore, both LIGO and ET have three detectors, but ET’s detectors are all coplanar while LIGO’s are not. The relative positions of LIGO’s three detectors offer greater advantages compared with ET. Despite having an SNR one order of magnitude higher than LIGO, ET still has very similar $\Delta\Omega$ values.
@@
Regarding the ground-based detector network, the currently operational LVK is unable to compensate for the sensitivity differences between second-generation and third-generation detectors. The limitations of ppE parameters in LVK differ by one to two orders of magnitude compared with ET. Furthermore, due to LVK’s varied response angles, $\Delta\Omega$ values are slightly superior to those of ET. The combination of LVK+ET provides the most comprehensive enhancement in network sensitivity. For instance, $\Delta\alpha_{Q}$ and $\Delta\alpha_{D}$ can approach the sensitivity level achieved by space-based detector networks, with potential surpassing in $\Delta\alpha_{V}$. Due to its arm length, the ground-based detector network’s $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ are four to six orders of magnitude higher than those observed in space-based detector networks, making it impossible to distinguish between breathing and longitudinal modes.
V.3 Results with luminosity distance
@@
The distances we assumed in our previous calculations and analysis were fixed: $D_{L}=44.6$ Mpc for SBBH and $D_{L}=6.79$ Gpc for MBHB. The GW signal can vary due to different distances, potentially affecting the final results. From Eq. (2), the strain is inversely proportional to the luminosity distance, represented as $h\propto 1/D_{L}$. Regarding the ppE parameter $\xi$, the partial derivative of the strain with respect to it, $\partial h/\partial\xi\propto 1/D_{L}$. Consequently, the FIM is $(\partial h/\partial\xi|\partial h/\partial\xi)\propto 1/D_{L}^{2}$, and thus the uncertainty of the ppE parameter scales as $\Delta\xi\propto D_{L}$. In other words, the uncertainty of the ppE parameter is directly proportional to the distance,
$\Delta\xi(D_{L})=\mathcal{K}D_{L}+\Delta\xi(0),$ (21)
@@
where $\mathcal{K}$ is a proportionality constant. The results for SBBH and MBHB are calculated at various distances, as presented in Table 2.
@@
$10^{5}$9.05 -12.63 -6.55 -4.86
LISA $10^{6}$8.76 -12.89 -6.46 -4.39
$10^{7}$9.50 -14.17 -6.64 -2.42
$10^{5}$9.35 -12.94 -6.83 -5.05
TaiJi $10^{6}$8.94 -13.11 -6.57 -4.71
$10^{7}$9.78 -14.46 -6.85 -2.68
$10^{5}$8.95 -12.58 -6.30 -3.12
TianQin $10^{6}$8.47 -12.73 -6.03 -1.72
$10^{7}$9.28 -13.99 -5.97 -0.25
$3$5.74 -9.54 -3.49 4.10
LIGO $20$6.25 -10.67 -4.02 3.57
$100$6.35 -10.38 -3.98 3.54
$3$5.41 -9.26 -1.27 6.25
Virgo $20$5.96 -10.38 -1.24 5.75
$100$6.03 -10.04 -0.96 5.35
$3$5.28 -9.14 -0.92 5.76
KAGRA $20$5.94 -10.40 -1.01 4.61
$100$5.73 -9.77 -0.58 3.97
$3$6.77 -10.45 -4.68 3.24
ET $20$7.20 -11.62 -5.13 2.79
$100$7.54 -11.64 -5.32 2.46
Equation (21) reveals that the uncertainty of the ppE parameters is directly proportional to the distance. The rate of change in uncertainty with distance varies for different ppE parameters. Table 2 illustrates that the influence of distance on $\Delta\alpha_{Q}$ and $\Delta\alpha_{D}$ is negligible, since most $\mathcal{K}$ values are less than $10^{-5}$. This can be attributed to the fact that these two parameters are closely linked to frequency, and any alteration in GW amplitude caused by distance has an negligible effect on frequency resolution. Moreover, ground-based detectors exhibit relatively larger $\mathcal{K}$ values compared with space-based detectors.
The $\mathcal{K}$ for $\Delta\alpha_{V}$ exhibits a significantly larger magnitude than the first two parameters. The $\mathcal{K}$ values of space-based detectors are typically around $10^{-6}$, while Virgo and KAGRA have $\mathcal{K}$ values around $10^{-1}$. LIGO and ET outperform Virgo and KAGRA, in most cases with $\mathcal{K}<10^{-3}$. These observations are consistent the analysis presented in the previous section, which elucidates these differences stem from detector performance.
Due to the degeneracy in ground-based detectors, $\Delta\alpha_{B}$ ($\Delta\alpha_{L}$) exhibits significantly large values of $\mathcal{K}$, exceeding $10^{2}$. In contrast, space-based detectors effectively resolve this degeneracy and improves outcomes. The results in Table 2 depend on the frequency range of MBHBs and the transfer frequency. The frequencies of smaller mass MBHBs are lower and closer to the transfer frequency, enhancing the resolution of these two modes and resulting in smaller $\mathcal{K}$ values. The results are consistent across different detectors: Taiji exhibits the lowest transfer frequency, whereas TianQin demonstrates the highest, aligning with their respective $\mathcal{K}$ values.
@@ V.4 Multiband observation
Space-based detectors can observe both MBHBs and SBBHs. Among numerous sources, only a few SBBHs rapidly increase in frequency and merge into the frequency band of ground-based detectors, enabling multiband observations by both space- and ground-based detectors [72]. The FIM is calculated for two types of SBBHs observed across multiple frequency bands and compared with results solely obtained from ground-based detectors, as shown in Fig. 11.
@@
Multiband observations greatly enhances the capabilities of ground-based detectors, particularly in effectively distinguishing between breathing and longitudinal modes. The addition of space-based detectors resolves the degeneracy, enabling multiband observations of SBBH to constrain $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ at levels comparable with those achieved by observing MBHB with space-based detectors. Additionally, the multiband observations greatly enhance the performance of Virgo and KAGRA in $\Delta\alpha_{V}$, reducing it by nearly three orders of magnitude. This brings them closer to the levels achieved by LIGO and ET. The $\Delta\alpha_{V}$ level of the ground-based detector network remains higher than that of the space-based detector network, even with multiband observations. Therefore, while there is improvement from these observations, its significance may not be as pronounced.
@@
The improvements for SBBH due to multiband observations vary with the mass, with respect to $\Delta\alpha_{Q}$ and $\Delta\alpha_{D}$. For SBBH with $M=20\ \mathrm{M_{\odot}}$, there is no significant improvement observed, while for a mass of $100\ \mathrm{M_{\odot}}$, some enhancement is demonstrated. The sensitivity of ET is so high that multiband observation does not significantly improve ET and LVK+ET, but it notably enhances the performance of only second-generation detectors. The frequency ranges of both types of SBBH differ within the space-based detector band, despite their one-year duration. The frequency variation is 46 mHz for SBBHs with $M=20\ \mathrm{M_{\odot}}$ and 78 mHz for those with $M=100\ \mathrm{M_{\odot}}$. The stronger constraints are caused by greater frequency variations, which in turn affect the observed improvements. Additionally, larger frequency variations occur with longer observation times. While four-year observations generally outperform one-year observations, this improvement is not significant. This is because the further away from the merger, the smaller the frequency change over the same time period. The SBBH with $M=100\ \mathrm{M_{\odot}}$ experiences a frequency shift of 78 mHz within one year before reaching 0.1 Hz. The change over four years was only 87 mHz, with the additional change within three years being just 9 mHz. Hence, our choice of a one-year observation duration remains justifiable.
@@
The performance of various space-based detectors significantly impacts the outcomes of multi-band observations. Moreover, the cutoff frequency of SBBHs within the space-based detector band is chosen to be 0.1 Hz, which is lower than TianQin’s transfer frequency of 0.28 Hz. That disparity renders TianQin less effective than LISA and Taiji in observing breathing and longitudinal modes, resulting in a much weaker multiband observation improvement for $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$. Conversely, TianQin exhibits significantly enhanced multiband observations for $\Delta\alpha_{Q}$, $\Delta\alpha_{D}$, and $\Delta\alpha_{V}$ due to its superior sensitivity within the selected SBBH frequency band compared with other detectors. The inclusion of space-based detectors compensates for the degeneracy observed in ground-based detectors regarding scalar modes and enhances other aspects. The advancement allows SBBH observation results to reach a level comparable to MBHB results.
@@ V.5 Multimessenger observation
Apart from using space- and ground-based detectors for multiband observations, integrating multimessenger observations, including electromagnetic (EM) observations, can also enhance GW detections. EM observations provide a distinctive viewpoint on BBH sources that is separate from GW observations. The studies conducted by Refs. [73, 74] have demonstrated that mergers of MBHBs can emit EM radiation from accretion disks. Additionally, the mergers of SBBHs in active galactic nuclei have the potential to generate powerful EM jets [73, 74, 75, 76]. These findings highlight the significance of both scenarios as promising targets for EM detectors. Accurate determination of the source’s sky position facilitates subsequent EM follow-up observations to identify counterparts. The findings from EM observations can then serve as valuable prior information to minimize uncertainties in GW observations.
@@
The EM effects generated by BBH mergers can be detected by infrared, optical, and x-ray observatories, with diverse implications for advancing GW research. For simplicity, we consider an ideal scenario where EM observations accurately determine specific parameters, disregarding discrepancies among EM detectors. According to Refs. [70, 47], the FIM removes the corresponding row and column when incorporating EM observations in order to mitigate uncertainties of other parameters in the GW data. This approach evaluates the performance enhancement of GW observations under ideal multimessenger conditions. The enhancements are quantified when EM observations accurately determine the distance, sky position, or inclination angle, and these results are compared with GW observations alone. Notably, although this study only focuse on ppE parameters, multimessenger observations also enhance time, mass, polarization, and other parameters. For details on enhancements in non-ppE parameters, please refer to Ref. [47].
@@
The multimessenger observations for LISA and Taiji demonstrate significant enhancements primarily for MBHBs with $M=10^{7}\ \mathrm{M_{\odot}}$. Utilizing multimessenger observations to determine the distance or inclination angle can result in a reduction of $\Delta\alpha_{V}$, $\Delta\alpha_{B}$, and $\Delta\alpha_{L}$ by 15% to 24%. The determination of the sky position can lead to a reduction in these parameters ranging from 36% to 49%. The multimessenger observations significantly enhance TianQin’s detection capabilities for all three typical MBHB masses. When accurately determining the distance, inclination angle, or sky position, $\Delta\alpha_{V}$ can be reduced by 40% to 57%. Furthermore, reductions in $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ are less, ranging from 9% to 17%.
@@
The impact of multimessenger observations on $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ is significantly diminished for ground-based detectors compared with multiband observations, as it fails to resolve the degeneracy and thus is not further considered. The improvement shown by LIGO is relatively less significant than that of ET, mainly because of its varied response angles. The determination of the distance or inclination angle leads to a reduction in $\Delta\alpha_{V}$ by 5% to 7%, while negligible changes are observed when determining the sky position. The enhancements exhibited by ET are more significant. When any one of the three parameters is determined, $\Delta\alpha_{V}$ can be reduced by 15% to 43%, with greater reductions observed for higher-mass SBBHs. Moreover, multimessenger observations significantly enhance Virgo and KAGRA, reducing $\Delta\alpha_{V}$ by 97% to 99.4%. The improvement achieved is only one order of magnitude lower than that observed in LIGO, demonstrating a significant advancement beyond GW observations alone.
@@
Multimessenger observations, whether for space- or ground-based detectors, do not improve the other two ppE parameters, $\Delta\alpha_{Q}$ and $\Delta\alpha_{D}$. The reason is that these two ppE parameters directly link to the highest frequency of GWs, where precision in observations is at its peak. Theoretically, multimessenger observations could enhance constraints on all ppE parameters, but in practice, some improvements only offer marginal benefits. The results presented in this study are limited to cases where the parameter uncertainty is reduced by more than 5%, with most unreported results reducing less than 1%. In conclusion, multimessenger observations can greatly improve the effectiveness of GW observations, resulting in stronger constraints on ppE parameters.